# WCCNet: Wavelet-context Cooperative Network for Efficient Multispectral Pedestrian Detection

### v2

@@ III Proposed Method
The overall architecture of WCCNet is shown in Fig. 2. The WCCNet mainly consists of three parts: (1) wavelet-integrated dual-stream backbone for fast and lightweight multispectral feature extraction; (2) crossmodal rearranging fusion module for semantic-aware crossmodal fusion; (3) decoupling detection head following YOLOX [32] for task-oriented predictions. In the following subsections, we shall mainly introduce the proposed backbone and fusion module.
III-A DWT-integrated Dual-stream Backbone
Overall Structure
For higher computational efficiency and precision, we propose a wavelet-integrated dual-stream backbone to extract features differentially according to the characteristics of different modalities. The proposed backbone consists of two branches: (1) RGB branch employs ordinary learnable convolutions to extract generic features from RGB modality in spatial domain; (2) Infrared branch utilizes wavelet-integrated layers to extract different frequency components for infrared images.
Adaptive DWT layers (ADWT) layer is composed of fixed wavelet kernels and learnable gain factors. The wavelet kernel is implemented by fixed-parameter and separable 1D convolution in two orthogonal directions, and the depth of infrared branch is quite shallower than RGB branch. In this way, the proposed backbone achieves fewer parameters and floating-point operations than ordinary symmetrical CNN backbones.
@@ III-A DWT-integrated Dual-stream Backbone
Let $\mathbf{X}_{R}\in\mathbb{R}^{3\times H\times W}$ and $\mathbf{X}_{T}\in\mathbb{R}^{1\times H\times W}$ denote the input tensors for RGB branch and infrared branch, where $H$ and $W$ represent the height and width of input tensors respectively. The output multiscale features of RGB branch and infrared branch at $j^{th}$ stage are defined as $\mathbf{I}_{R}^{j}\in\mathbb{R}^{C_{R}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ and $\mathbf{I}_{T}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$, where $j\in\{1,2,3\}$, $C_{R}^{j}$ and $C_{T}^{j}$ denote the channel numbers of these two branches at $j^{th}$ stage respectively. Additional two stages are applied to extract higher-level semantic features from the fusion of $\mathbf{X}_{R}$ and $\mathbf{X}_{T}$. Refined multispectral features from the last three stages are adopted for further pedestrian detection, namely $\mathbf{I}_{O}^{i}\in\mathbb{R}^{C^{i}_{O}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}}$, where $i\in\{3,4,5\}$ and $C^{i}_{O}$ denotes the channel numbers at $i^{th}$ stage. Following He et al. [37], residual blocks are employed in RGB branch and the last two stages.
Multiscale ADWT Layer
We now introduce infrared branch stacked by several multiscale ADWT layers in a sequence. Given is the input tensor $\mathbf{X}_{T}^{j-1}\in\mathbb{R}^{C_{TX}^{j-1}\times\frac{H}{2^{j-1}}\times\frac{W}{2^{j-1}}}$ from $(j-1)^{th}$ stage to $j^{th}$ stage. A set of 1D convolutions are performed by ADWT both vertically and horizontally to decompose $\mathbf{X}_{T}^{j-1}$ into a sequence of subdivided frequency components as the infrared feature $\mathbf{I}_{T}^{j}$. The obtained $\mathbf{I}_{T}^{j}$ is of double downsampled size, whose frequency contents vary in channel dimension. Then $\mathbf{I}_{T}^{j}$ will be subsequently fused with $\mathbf{I}_{R}^{j}$ to provide additional frequency-domain details.
@@
The ADWT layer can be regarded as the projection onto an orthogonal basis of small finite waves. Differing from learnable convolutions in CNN, the 1D convolutions conducted by ADWT have fixed weights and consist of two types of kernels, i.e., kernel $\mathbf{h}=(h_{0},h_{1},...,h_{K})$ as high-pass filter and kernel $\mathbf{g}=(g_{0},g_{1},...,g_{K})$ as low-pass filter, where $K$ denotes the kernel length. The corresponding weights for these filters are fixed and can be selected from different wavelet families.
@@
Several issues will arise for us to cope with. Wavelets with long support length are not able to separate close spaced features. Besides, unsymmetrical wavelets could lead to phase distortion in image processing. Therefore, symmetrical wavelets with short support length are chosen in this paper, specifically, Haar wavelet. Moreover, even-sized kernels of symmetrical wavelets will lead to shrinkage in size and potential shift problem. Thus, we symmetrically pad $\mathbf{X}_{T}^{j-1}$ following[38] before convolving $\mathbf{X}_{T}^{j-1}$ with wavelet kernels.
Then these 1D high-pass and low-pass filters convolve $\mathbf{X}_{T}^{j-1}$ by stride $s=2$ in both horizontal and vertical directions, resulting in approximation features $\mathbf{A}^{j}$, horizontal detailed features $\mathbf{DH}^{j}$ and vertical detailed features $\mathbf{DV}^{j}$. We assume that $\mathbf{d}^{j}_{:,m,n}$ and $\mathbf{a}^{j}_{:,m,n}$ are obtained by horizontal convolution with $\mathbf{h}$ and $\mathbf{g}$ on $\mathbf{X}_{T}^{j-1}$ respectively,
$\left\{\begin{aligned} \mathbf{d}^{j}_{:,m,n}&=\sum\nolimits_{k=1}\nolimits^{K}h_{k}\mathcal{P}(\mathbf{X}_{T}^{j-1})_{:,m,{sn+k}}\\ \mathbf{a}^{j}_{:,m,n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathcal{P}(\mathbf{X}_{T}^{j-1})_{:,m,{sn+k}}\end{aligned}\right$., (1)
@@
where $m$ and $n$ denote vertical and horizontal coordinates respectively. Note that $m\in\{0,1,...,\frac{H}{2^{j-1}}\}$ and $n\in\{0,1,...,\frac{W}{2^{j}}\}$. Function $\mathcal{P}(\cdot)$ represents symmetric padding in [38]. In the same way, $\mathbf{DH}^{j}$, $\mathbf{DV}^{j}$, and $\mathbf{A}^{j}$ can be obtained by vertical convolution, which are formulated as follows
$\left\{\begin{aligned} \mathbf{DH}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathbf{d}^{j}_{:,sm^{\prime}+k,n}\\ \mathbf{DV}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}h_{k}\mathbf{a}^{j}_{:,sm^{\prime}+k,n}\\ \mathbf{A}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathbf{a}^{j}_{:,sm^{\prime}+k,n}\end{aligned}\right$., (2)
where $m^{\prime}=(0,1,...,\frac{H}{2^{j}})$ and $n=(0,1,...,\frac{W}{2^{j}})$. Note that ADWT here does not include the highest frequency component $\mathbf{HH}^{j}$, which can be obtained by convolving with kernel $h$ both in horizontal and vertical directions. This choice comes from experimental results in Table IV of Sec. IV-D, which show that the network without $\mathbf{HH}^{j}$ achieves better performance. A possible explanation might be that the highest frequency component has little useful information for pediatrician detection, and its involvement might amplify underlying noise. As for other applications, it might be useful. Finally, we assign learnable gain factors to frequency components $\mathbf{DH}^{j}$, $\mathbf{DV}^{j}$, and $\mathbf{A}^{j}$, so that ADWT has frequency adaptability to datasets. The output tensor $\mathbf{I}_{T}^{j}$ at $j^{th}$ can be represented as follows,
$\mathbf{I}_{T}^{j}=\boldsymbol{\lambda}^{j}[\mathbf{DH}^{j},\mathbf{DV}^{j},\mathbf{A}^{j}],$ (3)
@@
where $[,]$ denotes concatenation at channel dimension. $\boldsymbol{\lambda}^{j}\in\mathbb{R}^{C_{T}^{j}}$ represents gain factors at $j^{th}$ stage adaptive to datasets, which is learned by depthwise $1\times 1$ convolution.
[…]
Analysis on Memory and Calculational Cost
To further analyze the savings in computational and memory cost of the proposed ADWT-integrated subnetwork, we consider a customized task of general settings as follows. Given input feature map $\mathbf{X}_{in}\in\mathbb{R}^{C^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ at $j^{th}$ stage with a spatial resolution of $\frac{H}{2^{j}}\times\frac{W}{2^{j}}$ and $C^{j}$ channels, the output feature map $\mathbf{I}_{out}\in\mathbb{R}^{C^{j+1}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$ is expected. Note that $C^{j}$ in RGB branch and infrared branch is different, while the former is much bigger than latter. Therefore, for fair efficiency comparison on the same task, the CE module introduced in Sec. III-B is also regarded as one part of the ADWT subnetwork in this analysis.
For the $j^{th}$ stage of CNN subnetwork, the first layer is a $K\times K$ CNN layer for down-sampling with a kernel $\theta_{0}\in\mathbb{R}^{C^{j}\times C^{j+1}\times K\times K}$ and a stride = 2. Then it is followed by a residual block. The residual block contains $2N$ CNN layers, in which $N$ CNN layers are of kernel size $K\times K$ and the other $N$ layers are of kernel size $1\times 1$. The $K\times K$ CNN layer in residual block is supposed to have a kernel $\theta_{1}\in\mathbb{R}^{C^{j+1}\times C^{j+1}\times K\times K}$ with a stride of 1 and a padding size of $\lfloor\frac{K-1}{2}\rfloor$. Following the computational complexity provided by He et al. [39] which ignores the computational cost of bias and batch normalization, the computational cost $P_{CNN}$ and the memory cost $M_{CNN}$ in the forward inference of CNN subnetwork at $j^{th}$ stage can be estimated as
$P_{CNN}\approx\frac{HW}{4^{j+1}}K^{2}\left(C^{j+1}\right)^{2}\left(\frac{C^{j}}{C^{j+1}}+\frac{N(K^{2}+1)}{K^{2}}\right)$ (4)
$M_{CNN}\approx K^{2}C^{j+1}C^{j}+N(K^{2}+1)\left(C^{j+1}\right)^{2}.$
Considering substitute CNN subnetwork with ADWT subnetwork and CE module at $j^{th}$ stage, the CE module introduced in Sec. III-B is also involved for fair comparison. The ADWT subnetwork first convolves the input feature map with paired low-pass and high-pass kernels $g\in\mathbb{R}^{1\times K\times 1}$ and $h\in\mathbb{R}^{1\times 1\times K}$ in vertical and horizontal directions. Then ADWT adopt depthwise $1\times 1$ convolution with learnable gain coefficients to obtain the weighted feature map $\mathbf{I}_{T}^{j}\in\mathbb{R}^{3C^{j}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$. Then $\mathbf{I}_{T}^{j}$ are convolved by the CE module in III-B with $K\times K$ convolutional layer to obtain the output feature map $\mathrm{I}_{out}$. Therefore, the computational cost $P_{DWT}$ and the memory cost $M_{DWT}$ in the forward propagation of ADWT subnetwork and CE module at $j^{th}$ stage are
$P_{ADWT+CE}\approx\frac{3HW}{4^{j+1}}C^{j}(\underbrace{K+1}_{ADWT}+\underbrace{2K^{2}C^{j+1}}_{CE})$ (5)
$M_{ADWT+CE}\approx\underbrace{2K+3C^{j}}_{ADWT}+\underbrace{6K^{2}C^{j}C^{j+1}}_{CE}.$
Moreover, since the weights of DWT layers are fixed and non-trainable, the ADWT subnetwork requires much lower computational cost in backward propagation, which means less training time. Given the gradient of the output feature map $\delta_{j+1}\in\mathbb{R}^{C^{j+1}\times\frac{H}{2^{j+1}}\times\frac{W}{2^{j+1}}}$, the gradient $\delta_{j}$ of the input feature map and the gradients $\gamma_{j}$ of the CNN kernels are expected. Assuming that there is no learnable bias in CNN layers, the computational cost in the backward propagation of CNN subnetwork at $j^{th}$ stage is
$B_{CNN}\approx\frac{HW}{4^{j}}C^{j+1}K^{2}(\frac{1}{4}NC^{j+1}+C^{j}).$ (6)
In the same way, the computational cost in the backward propagation of ADWT subnetwork at $j^{th}$ stage is
$B_{ADWT+CE}\approx\frac{HW}{4^{j}}(\underbrace{3C^{j}}_{ADWT}+\underbrace{\frac{1}{4}K^{2}C^{j}C^{j+1}}_{CE}).$ (7)
Therefore, in both forward and backward propagation, ADWT subnetwork has much lower computational complexity than the CNN subnetwork at the same stage, which also demonstrated by quantitative experiments in Sec. IV-D.
@@
Figure 3: Detailed Structure of CMRF, consisting of three parts: Cross-domain Embedding (CE) layer to narrow semantic gap, Crossmodal Spatial Alignment (CSA) layer to maintain spatial consistency, and Semantic Rearranging Fusion (SRF) layer to amplify crossmodal complementary information.
III-B Multi-domain Crossmodal Rearranging Fusion

### v3

@@ 3.1.1 Overall Structure
Figure 3: Overview framework of WCCNet. WCCNet mainly consists of three parts: (1) dual-stream backbone with MoWE and generic neural layers for fast and lightweight multispectral feature extraction; (2) CMRF module for semantic-aware crossmodal fusion; (3) decoupling detection head.
@@
Let $\mathbf{X}_{R}\in\mathbb{R}^{3\times H\times W}$ and $\mathbf{X}_{T}\in\mathbb{R}^{1\times H\times W}$ denote the input tensors for RGB branch and infrared branch, where $H$ and $W$ represent the height and width of input tensors respectively. The output multiscale features of RGB branch and infrared branch at $j^{th}$ stage are defined as $\mathbf{I}_{R}^{j}\in\mathbb{R}^{C_{R}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$ and $\mathbf{I}_{T}^{j}\in\mathbb{R}^{C_{T}^{j}\times\frac{H}{2^{j}}\times\frac{W}{2^{j}}}$, where $j\in\{1,2,3\}$, $C_{R}^{j}$ and $C_{T}^{j}$ denote the channel numbers of these two branches at $j^{th}$ stage respectively. Additional two stages are applied to extract higher-level semantic features from the fusion of $\mathbf{I}_{R}^{j}$ and $\mathbf{I}_{T}^{j}$. Refined multispectral features from the last three stages are adopted for further pedestrian detection, namely $\mathbf{I}_{O}^{i}\in\mathbb{R}^{C^{i}_{O}\times\frac{H}{2^{i}}\times\frac{W}{2^{i}}}$, where $i\in\{3,4,5\}$ and $C^{i}_{O}$ denotes the channel numbers at $i^{th}$ stage. Residual CNN blocks[he2016deep] are employed in RGB branch and the last two stages.
3.1.2 Mixture of Wavelet Experts
Infrared imagery generally exhibits weak textures but strong structural and thermal cues, requiring dynamic frequency-domain filtering. To capture informative infrared patterns with lower computational cost, we utilize Mixture of Wavelet Experts (MoWE) instead of ordinary heavyweight designs. Given input tensor $\mathbf{I}_{T}^{j-1}\in\mathbb{R}^{C_{T}^{j-1}\times\frac{H}{2^{j-1}}\times\frac{W}{2^{j-1}}}$ from $(j-1)^{th}$ stage to $j^{th}$ stage, MoWE leverages multispectral context-aware router to dynamically route different channels of $\mathbf{I}_{T}^{j-1}$ through sparsely activated ADWT experts and shared DWT experts. Then subdivided frequency patterns are captured as obtained infrared feature $\mathbf{I}_{T}^{j}$. $\mathbf{I}_{T}^{j}$ is of double downsampled size, and will be subsequently fused with $\mathbf{I}_{R}^{j}$ to provide complementary details.
Multispectral Context-aware Router. To assess environmental context jointly, multispectral context-aware router of MoWE first aggregates information from both current infrared input $\mathbf{I}_{T}^{j-1}$ and RGB feature $\mathbf{I}_{R}^{j}$. Given the number of routed experts $N_{e}$, routing score $\mathbf{S}_{e}^{j}\in\mathbb{R}^{c_{T}^{j}\times N_{e}}$ at $j$-th stage is obtained from multispectral context as follows,
$\mathbf{S}_{e}^{j}=\mathbf{W}\cdot\textrm{Flat}(\textrm{Cat}(\mathbb{P}(\mathbf{I}_{T}^{j-1}),\mathbb{P}(\mathbf{I}_{R}^{j})))\cdot(1+\lambda_{\epsilon}\epsilon),$ (1)
where $\textrm{Cat}(\cdot,\cdot)$ denotes channel concatenation, $\mathbb{P}(\cdot)$ represents adaptive average pooling operation to pool input into $H_{p}\times W_{p}$, and $\textrm{Flat}(\cdot)$ denotes flattening operation at last two channels. $\mathbf{W}\in\mathbb{R}^{N_{e}\times((C_{T}^{j-1}+C_{R}^{j})\cdot H_{p}\cdot W_{p})}$ is learnable weight of MLP layer. $\epsilon$ is standard Gaussian noise with hyperparameter $\lambda_{\epsilon}$ to control noise level. $\epsilon$ enable MoWE to explore diverse routing strategies during training, and is removed during inference. $\lambda_{\epsilon}$ is set to 0.2 in this paper. Then activated experts are selected based on top-$N_{act}$ scores of $\mathbf{S}_{e}^{j}$ for each channel, resulting in activating score $\mathbf{S}_{\Omega}^{j}\in\mathbb{R}^{C_{T}^{j-1}\times N_{act}}$, where $N_{act}$ is the number of activated experts for each channel, as follows,
$\mathbf{S}_{\Omega}^{j}=\operatorname{Softmax}(\{\mathbf{S}_{e}^{j}[:,k]\mid k\in\Omega\}),\;\Omega=\operatorname{Top-K}(\mathbf{S}_{e}^{j}),$ (2)
where $\operatorname{Top-K}(\cdot)$ denotes top-$N_{act}$ selection operation from $N_{e}$ scores, and $\omega$ is the activated set of experts for each channel. $\mathbf{S}_{e}^{j}[:,k]$ means $k^{th}$ column of $\mathbf{S}_{e}^{j}$. $\mathbf{S}_{\Omega}^{j}$ and $\Omega$ are then utilized in subsequent expert aggregation.
@@ 3.1.2 Mixture of Wavelet Experts
Heterogeneous Wavelet Experts. Learnable ADWT and static DWT are involved in MoWE. They can be regarded as the projection onto an orthogonal basis of small finite waves, and both perform 1D convolutions with two types of kernels, i.e., kernel $\mathbf{h}=(h_{0},h_{1},...,h_{K})$ as high-pass filter and kernel $\mathbf{g}=(g_{0},g_{1},...,g_{K})$ as low-pass filter. $K$ denotes kernel length. The corresponding weights for these filters are initialized from pre-defined wavelet families. Notably, there are two differences between ADWT and DWT. Firstly, filter kernels $\mathbf{h}$ and $\mathbf{g}$ of ADWT are learnable, and $\mathbf{h}$ is constrained to conjugate mirror filter of $\mathbf{g}$ following [PNASWavelet]. Secondly, the extracted frequency components of ADWT are further modulated by gain factors.
@@
Besides, for wavelet transform on infrared features, there exist several issues for us to cope with. Wavelets with long support length are not able to separate close spaced features. Besides, unsymmetrical wavelets could lead to phase distortion in image processing. Therefore, symmetrical wavelets with short support length are chosen in this paper. Moreover, even-sized kernels of symmetrical wavelets will lead to shrinkage in size and potential shift problem. Thus, we symmetrically pad $\mathbf{I}_{T}^{j-1}$ following[wu2019convolution] before convolving $\mathbf{I}_{T}^{j-1}$ with wavelet kernels. Then these 1D high-pass and low-pass filters convolve $\mathbf{I}_{T}^{j-1}$ by stride $s=2$ in both horizontal and vertical directions, resulting in approximation features $\mathbf{A}^{j}$, horizontal detailed features $\mathbf{DH}^{j}$, vertical detailed features $\mathbf{DV}^{j}$, and diagonal detailed $\mathbf{HH}^{j}$. We assume that $\mathbf{d}^{j}_{:,m,n}$ and $\mathbf{a}^{j}_{:,m,n}$ are obtained by horizontal convolution with $\mathbf{h}$ and $\mathbf{g}$ on $\mathbf{I}_{T}^{j-1}$ respectively,
$\left\{\begin{aligned} \mathbf{d}^{j}_{:,m,n}&=\sum\nolimits_{k=1}\nolimits^{K}h_{k}\mathcal{P}(\mathbf{I}_{T}^{j-1})_{:,m,{sn+k}}\\ \mathbf{a}^{j}_{:,m,n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathcal{P}(\mathbf{I}_{T}^{j-1})_{:,m,{sn+k}}\end{aligned}\right$., (3)
@@
where $m$ and $n$ denote vertical and horizontal coordinates respectively. Note that $m\in\{0,1,...,\frac{H}{2^{j-1}}\}$ and $n\in\{0,1,...,\frac{W}{2^{j}}\}$. Function $\mathcal{P}(\cdot)$ represents symmetric padding in [wu2019convolution]. In the same way, infrared frequency features can be obtained by vertical convolution,
$\left\{\begin{aligned} \mathbf{HH}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}h_{k}\mathbf{d}^{j}_{:,sm^{\prime}+k,n}\\ \mathbf{DH}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathbf{d}^{j}_{:,sm^{\prime}+k,n}\\ \mathbf{DV}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}h_{k}\mathbf{a}^{j}_{:,sm^{\prime}+k,n}\\ \mathbf{A}^{j}_{:,m^{\prime},n}&=\sum\nolimits_{k=1}\nolimits^{K}g_{k}\mathbf{a}^{j}_{:,sm^{\prime}+k,n}\end{aligned}\right$., (4)
where $m^{\prime}=(0,1,...,\frac{H}{2^{j}})$ and $n=(0,1,...,\frac{W}{2^{j}})$. Finally, for ADWT, we assign learnable gain factors to modulate frequency components to enhance frequency adaptability to pedestrian detection. The output of $i$-th ADWT and DWT experts at $j^{th}$ can be respectively represented as follows,
$\left\{\begin{aligned} \mathcal{E}_{act}[i](I_{T}^{j-1})&=\boldsymbol{\lambda}^{j}\cdot\textrm{Cat}(\mathbf{A}^{j},\mathbf{DH}^{j},\mathbf{DV}^{j},\mathbf{HH}^{j}),\\ \mathcal{E}_{shared}[i](I_{T}^{j-1})&=\textrm{Cat}(\mathbf{A}^{j},\mathbf{DH}^{j},\mathbf{DV}^{j},\mathbf{HH}^{j})\end{aligned}\right$. (5)
@@
where $\textrm{Cat}(\cdot,\cdot)$ denotes concatenation at channel dimension. $\boldsymbol{\lambda}^{j}\in\mathbb{R}^{C_{T}^{j}}$ represents gain factors at $j^{th}$ stage adaptive to datasets, learned by depthwise $1\times 1$ convolution.
Expert Aggregation. Given $N_{s}$ shared experts $\mathcal{E}_{shared}$ and $N_{act}$ activated experts $\mathcal{E}_{act}$ of $N_{e}$ routed experts, the output infrared feature $\mathbf{I}_{T}^{j}$ at $j^{th}$ stage is as follows,
$\mathbf{I}_{T}^{j}=\sum_{i\in\Omega}\textrm{Re}(\mathbf{S}_{\Omega}^{j}[:,i],4)\cdot\mathcal{E}_{act}[i](\mathbf{X}_{T}^{j-1})+\sum_{i=0}^{N_{s}}\mathcal{E}_{shared}(\mathbf{X}_{T}^{j-1}),$ (6)
where $\mathbf{S}_{\Omega}^{j}[:,i]$ means $i^{th}$ column of $\mathbf{S}_{\Omega}^{j}$. $\textrm{Re}(\cdot,4)$ operation repeats the first dimension of $\mathbf{S}_{\Omega}^{j}[:,i]$ for 4 times. $\mathbf{I}_{T}^{j}$ is then fused with RGB feature $\mathbf{I}_{R}^{j}$ by CMRF module to supplement crossmodal complementary information, and transferred to next stage for further feature extraction.
Channel-wise Routing Balance Loss. To prevent routing collapse of MoWE during training, following prior work [MoE], we apply channel-wise routing balance loss $\mathcal{L}_{b}$ to the routing scores $\mathbf{S}_{e}^{j}$ at three stages. Let $\eta_{k,j}=\sum_{c=1}^{C_{T}^{j}}\mathds{1}(k\in\Omega[c])$ be routing frequency of $k$-th expert across channel $c$, where $\mathds{1}(\cdot)$ denotes the indicator function and $\Omega[c]$ denotes the subset $\operatorname{Top-K}(\mathbf{S}_{e}^{j}[c])$ in Eq. 2. And let $P_{k,j}=\sum_{c=1}^{C_{T}^{j}}\operatorname{Softmax}(\mathbf{S}_{e}^{j})[c,k]$ be the average probability of $k$-th expert assigned by the router. $\mathcal{L}_{b}$ is defined as
$\mathcal{L}_{b}=\sum_{j=1}^{3}\frac{N_{e}}{(C_{T}^{j})^{2}}\sum_{k=1}^{N_{e}}\eta_{k,j}\cdot P_{k,j}.$ (7)
$\mathcal{L}_{b}$ is added to the total loss with a weight of 0.1 in practice.
@@
Figure 4: Detailed structure of CMRF. It consists of three parts: Cross-domain Embedding (CE) layer to narrow semantic gap, Crossmodal Spatial Alignment (CSA) layer to maintain spatial consistency, and Semantic Rearranging Fusion (SRF) layer to amplify crossmodal complementary information.
3.2 Multi-domain Crossmodal Rearranging Fusion